# Bootstrapping Estimators based on the Block Maxima Method

### v1

@@ 11.1 A general result on bootstrapping the Fréchet independence MLE
$\lim_{n\to\infty}\mathbb{P}(X_{n,1}=\dots=X_{n,\omega_{n}})=0.$ (11.4)
For some $\alpha_{0}>0$ and $\sigma_{n}\to\infty$, proximity of the $F_{n,1},\dots,F_{n,\omega_{n}}$ to the Fréchet-distribution $P_{\alpha_{0},\sigma_{n}}$ will be controlled by convergence conditions on empirical means of $X_{n,i}/\sigma_{n}$: for $0<\alpha_{-}<\alpha_{0}<\alpha_{+}<\infty$, let $\mathcal{F}(\alpha_{-},\alpha_{+})$ denote the set of functions containing $x\mapsto\log x,x\mapsto x^{-\alpha},x\mapsto x^{-\alpha}\log x$ and $x\mapsto x^{-\alpha}\log^{2}x$ for all $\alpha\in(\alpha_{-},\alpha_{+})$. Moreover, recall $f_{1}(x)=x^{-\alpha_{0}}$, $f_{2}(x)=x^{-\alpha_{0}}\log x$ and $f_{3}(x)=\log x$ from (5.4), considered as functions on $(0,\infty)$, and let =$\mathcal{H}=\{f_{1},f_{2},f_{3}\}$. For a real-valued function $f$ defined on $(0,\infty)$ such that the following integrals/expectations exist, let
@@
Recall the log-likelihood function $\ell_{\bm{\theta}}$ of the Fréchet distribution from (5.2).
Theorem 11.6 (Bücher and Segers, 2018b, ).
(i) Suppose that (11.4) holds and that Condition 11.5(i) is met. Then, on the complement of the event $\{X_{n,1}=\dots=X_{n,\omega_{n}}\}$, the independence Fréchet log-likelihood ${\bm{\theta}}\mapsto\sum_{i=1}^{\omega_{n}}\ell_{\bm{\theta}}(X_{n,i})$ has a unit maximizer $\hat{\bm{\theta}}_{n}=(\hat{\alpha}_{n},\hat{\sigma}_{n})$, and that maximizer is consistent in the sense that $(\hat{\alpha}_{n},\hat{\sigma}_{n}/\sigma_{n})=(\alpha_{0},1)+o_{\mathbb{P}}(1)$ as $n\to\infty$.
(ii) If, additionally, Condition 11.5(ii)-(iii) is met, then
@@
$\lim_{n\to\infty}\mathbb{P}(X_{n,1}^{*}=\dots=X_{n,\omega_{n}}^{*})=0$ (11.5)
@@
Theorem 11.7.
Suppose that (11.4), (11.5) and Condition 11.5 is met, that ^$\hat{\mathbb{P}}_{n}^{*}f=P_{\alpha_{0},1}f+o_{\mathbb{P}}(1)$ for all $f\in\mathcal{F}(\alpha_{-},\alpha_{+})$ (or, equivalently, (|^$\mathbb{P}(|\hat{\mathbb{P}}_{n}^{*}f-P_{\alpha_{0},1}|>\varepsilon\mid\mathcal{X}_{n})=o_{\mathbb{P}}(1)$ for all $\varepsilon>0$ and all $f\in\mathcal{F}(\alpha_{-},\alpha_{+})$) and that
$d_{K}\big{(}\mathcal{L}(\hat{\mathbb{G}}_{n}^{\ast}(f_{1},f_{2},f_{3})^{\top}\mid\mathcal{X}_{n}),\mathcal{N}_{3}(\bm{0},\Sigma_{G})\big{)}=o_{\mathbb{P}}(1).$ (11.6)
@@
$v_{n}\begin{pmatrix}\hat{\alpha}_{n}^{\ast}-\hat{\alpha}_{n}\\ \hat{\sigma}_{n}^{*}/\hat{\sigma}_{n}-1\end{pmatrix}=M(\alpha_{0})\hat{\mathbb{G}}_{n}^{*}(f_{1},f_{2},f_{3})^{\top}+o_{\mathbb{P}}(1).$ (11.7)
@@
$d_{K}\left[\mathcal{L}\left(v_{n}\begin{pmatrix}\hat{\alpha}_{n}^{\ast}-\hat{\alpha}_{n}\\ \hat{\sigma}_{n}^{*}/\hat{\sigma}_{n}-1\end{pmatrix}\,\Big{|}\,\mathcal{X}_{n}\right),\mathcal{L}\left(v_{n}\begin{pmatrix}\hat{\alpha}_{n}-\alpha_{0}\\ \hat{\sigma}_{n}/\sigma_{n}-1\end{pmatrix}\right)\right]=o_{\mathbb{P}}(1).$ (11.8)
Since it concerns empirical means only, the condition in (11.6) is typically a standard result for the bootstrap; see for instance Section 10 and 11 in Kosorok, (2008).
Proof of Theorem 11.7.
First, by Theorem 11.6 (which is Theorem 2.3, Theorem 2.5 and Addendum 2.6 in Bücher and Segers, 2018b, applied to the sample $\mathcal{X}_{n}$), we have
$v_{n}\begin{pmatrix}\hat{\alpha}_{n}-\alpha_{0}\\ \hat{\sigma}_{n}/\sigma_{n}-1\end{pmatrix}=M(\alpha_{0})(\bar{\mathbb{G}}_{n}+B_{n})(f_{1},f_{2},f_{3})^{\top}+o_{\mathbb{P}}(1)\rightsquigarrow M(\alpha_{0})\mathcal{N}_{3}(\bm{B}_{G},\Sigma_{G}).$ (11.9)
Next, in view of the fact that (11.6) implies $\hat{\mathbb{G}}_{n}^{*}(f_{1},f_{2},f_{3})^{\top}\rightsquigarrow\mathcal{N}_{3}(\bm{0},\Sigma_{G})$ (unconditionally) by Lemma 2.3 in Bücher and Kojadinovic, (2019), we have
@@
Hence, we may apply Theorem 2.3, Theorem 2.5 and Addendum 2.6 in Bücher and Segers, 2018b to the sample $\mathcal{X}_{n}^{*}$ (unconditionally) to obtain that
$v_{n}\begin{pmatrix}\hat{\alpha}_{n}^{\ast}-\alpha_{0}\\ \hat{\sigma}_{n}^{\ast}/\sigma_{n}-1\end{pmatrix}=M(\alpha_{0}){\mathbb{G}}_{n}^{\circ}(f_{1},f_{2},f_{3})^{\top}+o_{\mathbb{P}}(1)$ (11.10)
As a consequence, by (11.9), (11.10) and $\hat{\sigma}_{n}/\sigma_{n}=1+o_{\mathbb{P}}(1)$, we have
@@
$=M(\alpha_{0})\hat{\mathbb{G}}_{n}^{\ast}(f_{1},f_{2},f_{3})^{\top}+o_{\mathbb{P}}(1),$ (11.11)
which is (11.7).
Finally, for the proof of (11.8), let $\mathcal{X}_{n}^{\#}$ denote a second bootstrap sample, generated in the same way as $\mathcal{X}_{n}^{*}$ and independent of $\mathcal{X}_{n}^{*}$, conditionally on $\mathcal{X}_{n}$. Denote the respective estimators and empirical measures/processes by $\hat{\alpha}_{n}^{\#},\hat{\mathbb{P}}_{n}^{\#}$ etc. Then, by the expansions in (11.1),
@@
Equation (11.6) and Lemma 2.3 in Bücher and Kojadinovic, (2019) implies that the dominating term on the right-hand side converges weakly (unconditionally) to $\mathcal{N}_{2}(\bm{0},M(\alpha_{0})\Sigma_{G}M(\alpha_{0})^{\top})^{\otimes 2}$, which by another reverse application of that lemma implies
@@
The assertion then follows from the triangular inequality and (11.9), noting that $\bm{B}_{G}=\bm{0}$ by assumption. ∎
11.2 Auxiliary results on bootstrap consistency
Lemma 11.8.
@@ 11.2 Auxiliary results on bootstrap consistency
$d_{K}\big{[}\mathcal{L}(\bm{A}_{n}\bm{S}_{n}),\mathcal{L}(\bm{A}_{n}\bm{T}_{n})\big{]}=d_{K}\big{[}\mathcal{L}(\tilde{\bm{A}}_{n}\bm{S}_{n}),\mathcal{L}(\tilde{\bm{A}}_{n}\bm{T}_{n})\big{]}.$ (11.12)
Since $\tilde{\bm{A}}_{n}\in[-1,1]^{p\times q}$, the Bolzano-Weierstraß Theorem allows to find a subsequence $n^{\prime}:=n^{\prime}(n)$ such that $\bm{E}=\lim_{n\to\infty}\tilde{\bm{A}}_{n^{\prime}}$ exists. Slutsky’s lemma then yields the weak convergences $\tilde{\bm{A}}_{n^{\prime}}\bm{S}_{n^{\prime}}\rightsquigarrow\bm{E}\bm{S}$ and $\tilde{\bm{A}}_{n^{\prime}}\bm{T}_{n^{\prime}}\rightsquigarrow\bm{E}\bm{S}$. We will show below that the cdf of $\bm{E}\bm{S}$ is continuous. Since the Kolmogorov-metric $d_{K}$ metrizes weak convergence to limits with continuous cdfs (van der Vaart, , 1998, Lemma 2.11), an application of the triangular inequality implies
@@
for $n^{\prime}\to\infty$. As the previous argumentation can be repeated for subsequences $n^{\prime}$ of arbitrary subsequences $n^{\prime\prime},$ we then have $d_{K}\big{[}\mathcal{L}(\tilde{\bm{A}}_{n}\bm{S}_{n}),\mathcal{L}(\tilde{\bm{A}}_{n}\bm{T}_{n})\big{]}=o(1)$, which implies the assertion by (11.12).
@@
Hence, since $\bm{S}_{n}\rightsquigarrow Q$ by assumption, we obtain $\mathcal{L}(\bm{S}_{n^{\prime\prime}}^{\ast}\mid\bm{X}_{n^{\prime\prime}})\rightsquigarrow Q$ on $\Omega_{0}.$ Lemma 11.8 then implies

### v2

@@ C.1 A general result on bootstrapping the Fréchet independence MLE
$\lim_{n\to\infty}\mathbb{P}(X_{n,1}=\dots=X_{n,\omega_{n}})=0.$ (C.4)
For some $\alpha_{0}>0$ and $\sigma_{n}\to\infty$, proximity of the $F_{n,1},\dots,F_{n,\omega_{n}}$ to the Fréchet-distribution $P_{\alpha_{0},\sigma_{n}}$ will be controlled by convergence conditions on empirical means of $X_{n,i}/\sigma_{n}$: for $0<\alpha_{-}<\alpha_{0}<\alpha_{+}<\infty$, let $\mathcal{F}(\alpha_{-},\alpha_{+})$ denote the set of functions containing $x\mapsto\log x,x\mapsto x^{-\alpha},x\mapsto x^{-\alpha}\log x$ and $x\mapsto x^{-\alpha}\log^{2}x$ for all $\alpha\in(\alpha_{-},\alpha_{+})$. Moreover, recall $f_{1}(x)=x^{-\alpha_{0}}$, $f_{2}(x)=x^{-\alpha_{0}}\log x$ and $f_{3}(x)=\log x$ from (A.4), considered as functions on $(0,\infty)$, and let =$\mathcal{H}=\{f_{1},f_{2},f_{3}\}$. For a real-valued function $f$ defined on $(0,\infty)$ such that the following integrals/expectations exist, let
@@
Recall the log-likelihood function $\ell_{\bm{\theta}}$ of the Fréchet distribution from (A.2).
Theorem C.6 (BucSeg18a).
(i) Suppose that (C.4) holds and that Condition C.5(i) is met. Then, on the complement of the event $\{X_{n,1}=\dots=X_{n,\omega_{n}}\}$, the independence Fréchet log-likelihood ${\bm{\theta}}\mapsto\sum_{i=1}^{\omega_{n}}\ell_{\bm{\theta}}(X_{n,i})$ has a unit maximizer $\hat{\bm{\theta}}_{n}=(\hat{\alpha}_{n},\hat{\sigma}_{n})$, and that maximizer is consistent in the sense that $(\hat{\alpha}_{n},\hat{\sigma}_{n}/\sigma_{n})=(\alpha_{0},1)+o_{\mathbb{P}}(1)$ as $n\to\infty$.
(ii) If, additionally, Condition C.5(ii)-(iii) is met, then
@@
$\lim_{n\to\infty}\mathbb{P}(X_{n,1}^{*}=\dots=X_{n,\omega_{n}}^{*})=0$ (C.5)
@@
Theorem C.7.
Suppose that (C.4), (C.5) and Condition C.5 is met, that ^$\hat{\mathbb{P}}_{n}^{*}f=P_{\alpha_{0},1}f+o_{\mathbb{P}}(1)$ for all $f\in\mathcal{F}(\alpha_{-},\alpha_{+})$ (or, equivalently, (|^$\mathbb{P}(|\hat{\mathbb{P}}_{n}^{*}f-P_{\alpha_{0},1}|>\varepsilon\mid\mathcal{X}_{n})=o_{\mathbb{P}}(1)$ for all $\varepsilon>0$ and all $f\in\mathcal{F}(\alpha_{-},\alpha_{+})$) and that
$d_{K}\big{(}\mathcal{L}(\hat{\mathbb{G}}_{n}^{\ast}(f_{1},f_{2},f_{3})^{\top}\mid\mathcal{X}_{n}),\mathcal{N}_{3}(\bm{0},\Sigma_{G})\big{)}=o_{\mathbb{P}}(1).$ (C.6)
@@
$v_{n}\begin{pmatrix}\hat{\alpha}_{n}^{\ast}-\hat{\alpha}_{n}\\ \hat{\sigma}_{n}^{*}/\hat{\sigma}_{n}-1\end{pmatrix}=M(\alpha_{0})\hat{\mathbb{G}}_{n}^{*}(f_{1},f_{2},f_{3})^{\top}+o_{\mathbb{P}}(1).$ (C.7)
@@
$d_{K}\left[\mathcal{L}\left(v_{n}\begin{pmatrix}\hat{\alpha}_{n}^{\ast}-\hat{\alpha}_{n}\\ \hat{\sigma}_{n}^{*}/\hat{\sigma}_{n}-1\end{pmatrix}\,\Big{|}\,\mathcal{X}_{n}\right),\mathcal{L}\left(v_{n}\begin{pmatrix}\hat{\alpha}_{n}-\alpha_{0}\\ \hat{\sigma}_{n}/\sigma_{n}-1\end{pmatrix}\right)\right]=o_{\mathbb{P}}(1).$ (C.8)
Since it concerns empirical means only, the condition in (C.6) is typically a standard result for the bootstrap; see for instance Section 10 and 11 in Kos08.
Proof of Theorem C.7.
First, by Theorem C.6 (which is Theorem 2.3, Theorem 2.5 and Addendum 2.6 in BucSeg18a applied to the sample $\mathcal{X}_{n}$), we have
$v_{n}\begin{pmatrix}\hat{\alpha}_{n}-\alpha_{0}\\ \hat{\sigma}_{n}/\sigma_{n}-1\end{pmatrix}=M(\alpha_{0})(\bar{\mathbb{G}}_{n}+B_{n})(f_{1},f_{2},f_{3})^{\top}+o_{\mathbb{P}}(1)\rightsquigarrow M(\alpha_{0})\mathcal{N}_{3}(\bm{B}_{G},\Sigma_{G}).$ (C.9)
Next, in view of the fact that (C.6) implies $\hat{\mathbb{G}}_{n}^{*}(f_{1},f_{2},f_{3})^{\top}\rightsquigarrow\mathcal{N}_{3}(\bm{0},\Sigma_{G})$ (unconditionally) by Lemma 2.3 in BucKoj2019note, we have
@@
Hence, we may apply Theorem 2.3, Theorem 2.5 and Addendum 2.6 in BucSeg18a to the sample $\mathcal{X}_{n}^{*}$ (unconditionally) to obtain that
$v_{n}\begin{pmatrix}\hat{\alpha}_{n}^{\ast}-\alpha_{0}\\ \hat{\sigma}_{n}^{\ast}/\sigma_{n}-1\end{pmatrix}=M(\alpha_{0}){\mathbb{G}}_{n}^{\circ}(f_{1},f_{2},f_{3})^{\top}+o_{\mathbb{P}}(1)$ (C.10)
As a consequence, by (C.9), (C.10) and $\hat{\sigma}_{n}/\sigma_{n}=1+o_{\mathbb{P}}(1)$, we have
@@
$=M(\alpha_{0})\hat{\mathbb{G}}_{n}^{\ast}(f_{1},f_{2},f_{3})^{\top}+o_{\mathbb{P}}(1),$ (C.11)
which is (C.7).
Finally, for the proof of (C.8), let $\mathcal{X}_{n}^{\#}$ denote a second bootstrap sample, generated in the same way as $\mathcal{X}_{n}^{*}$ and independent of $\mathcal{X}_{n}^{*}$, conditionally on $\mathcal{X}_{n}$. Denote the respective estimators and empirical measures/processes by $\hat{\alpha}_{n}^{\#},\hat{\mathbb{P}}_{n}^{\#}$ etc. Then, by the expansions in (C.1),
@@
Equation (C.6) and Lemma 2.3 in BucKoj2019note implies that the dominating term on the right-hand side converges weakly (unconditionally) to $\mathcal{N}_{2}(\bm{0},M(\alpha_{0})\Sigma_{G}M(\alpha_{0})^{\top})^{\otimes 2}$, which by another reverse application of that lemma implies
@@
The assertion then follows from the triangular inequality and (C.9), noting that $\bm{B}_{G}=\bm{0}$ by assumption. ∎
C.2 Auxiliary results on bootstrap consistency
Lemma C.8.
@@ C.2 Auxiliary results on bootstrap consistency
$d_{K}\big{[}\mathcal{L}(\bm{A}_{n}\bm{S}_{n}),\mathcal{L}(\bm{A}_{n}\bm{T}_{n})\big{]}=d_{K}\big{[}\mathcal{L}(\tilde{\bm{A}}_{n}\bm{S}_{n}),\mathcal{L}(\tilde{\bm{A}}_{n}\bm{T}_{n})\big{]}.$ (C.12)
Since $\tilde{\bm{A}}_{n}\in[-1,1]^{p\times q}$, the Bolzano-Weierstraß Theorem allows to find a subsequence $n^{\prime}:=n^{\prime}(n)$ such that $\bm{E}=\lim_{n\to\infty}\tilde{\bm{A}}_{n^{\prime}}$ exists. Slutsky’s lemma then yields the weak convergences $\tilde{\bm{A}}_{n^{\prime}}\bm{S}_{n^{\prime}}\rightsquigarrow\bm{E}\bm{S}$ and $\tilde{\bm{A}}_{n^{\prime}}\bm{T}_{n^{\prime}}\rightsquigarrow\bm{E}\bm{S}$. We will show below that the CDF of $\bm{E}\bm{S}$ is continuous. Since the Kolmogorov-metric $d_{K}$ metrizes weak convergence to limits with continuous CDFs (Van98, Lemma 2.11), an application of the triangular inequality implies
@@
for $n^{\prime}\to\infty$. As the previous argumentation can be repeated for subsequences $n^{\prime}$ of arbitrary subsequences $n^{\prime\prime},$ we then have $d_{K}\big{[}\mathcal{L}(\tilde{\bm{A}}_{n}\bm{S}_{n}),\mathcal{L}(\tilde{\bm{A}}_{n}\bm{T}_{n})\big{]}=o(1)$, which implies the assertion by (C.12).
@@
Hence, since $\bm{S}_{n}\rightsquigarrow Q$ by assumption, we obtain $\mathcal{L}(\bm{S}_{n^{\prime\prime}}^{\ast}\mid\bm{X}_{n^{\prime\prime}})\rightsquigarrow Q$ on $\Omega_{0}.$ Lemma C.8 then implies